# Efficient Tactile Perception with Soft Electrical Impedance Tomography and Pre-trained Transformer

### v1

@@ VII Conclusion
This work introduces the PTET model, a novel framework for EIT-based tactile reconstruction that achieves high-resolution tactile mapping with minimal labelled real-world data through self-supervised pretraining. PTET mitigates the reliance on extensive annotated datasets and requires only 0.56% of the annotated samples compared to the SOTA DNN model (2,500 vs. 450,000) while delivering superior tactile reconstructions. Compared to its supervised counterpart, PTET-SL, PTET achieves comparable performance using just 5% of the annotated samples (2,500 vs. 50,000). Experimental results further demonstrate PTET’s ability to bridge the simulation-to-reality gap and surpass the SOTA DNN model in reconstruction accuracy, achieving up to a 63.88% improvement in overall performance. These findings validate PTET as an effective and scalable solution for tactile perception systems, paving the way for advanced applications in HMI and robotics. Beyond tactile sensing, the pre-trained encoder, based on the EIM representation, can be easily applied to other EIT sensing tasks or serve as a teacher model for developing a versatile EIT foundation model.

### v2

@@ VII Conclusion
This work introduces the PTET model, a novel framework for EIT-based tactile reconstruction that achieves high-resolution tactile mapping with minimal labelled real-world data through self-supervised pretraining. PTET mitigates the reliance on extensive annotated datasets and requires only 0.56% of the annotated samples compared to the SOTA DNN model (2,500 vs. 450,000) while delivering superior tactile reconstructions. Compared to its supervised counterpart, PTET-SL, PTET achieves comparable performance using just 5% of the annotated samples (2,500 vs. 50,000). This is primarily due to our lightweight linear projection strategy during fine-tuning; in contrast, heavier models with more complex mapping strategies are difficult to train effectively with only 2,500 paired samples (see Supplementary Material). Experimental results further demonstrate PTET’s ability to bridge the simulation-to-reality gap and surpass the SOTA DNN model in reconstruction accuracy, achieving up to a 63.88% improvement in overall performance. These findings validate PTET as an effective and scalable solution for tactile perception systems, paving the way for advanced applications in HMI and robotics. In future work, we plan to explore sensor miniaturization to enable high-resolution tactile sensing in compact or embedded platforms, further expanding the application scope of PTET in wearable, prosthetic, and soft robotic systems. Beyond tactile sensing, the pre-trained encoder, based on the EIM representation, can be easily applied to other EIT sensing tasks or serve as a teacher model for developing a versatile EIT foundation model. Pretrained models with few-shot fine-tuning reduce the need for large labeled datasets in industrial measurements, enabling rapid adaptation to new conditions with minimal data. By capturing transferable features, they enhance robustness and enable cost-effective, precise, and reliable measurement with reduced downtime and calibration effort.